# Navigating the phase diagram of quantum many-body systems in phase space

### v1

@@ Abstract
We demonstrate the unique capabilities of the Wigner function, particularly in its positive and negative parts, for exploring the phase diagram of the spin$-(\frac{1}{2\!}-\!\frac{1}{2})$ and spin$-(\frac{1}{2}\!-\!1)$ Ising-Heisenberg chains. We highlight the advantages and limitations of the phase space approach in comparison with the entanglement concurrence in detecting phase boundaries. We establish that the equal angle slice approximation in the phase space is an effective method for capturing the essential features of the phase diagram, but falls short in accurately assessing the negativity of the Wigner function for the homogeneous spin$-(\frac{1}{2}\!-\!\frac{1}{2})$ Ising-Heisenberg chain. In contrast, we find for the inhomogeneous spin$-(\frac{1}{2}\!-\!1)$ chain that an integral over the entire phase space is necessary to accurately capture the phase diagram of the system. This distinction underscores the sensitivity of phase space methods to the homogeneity of the quantum system under consideration.
@@ I Introduction
The rapid development of quantum technologies in the 21st century allowed for the possibility of experimental measurement of the phase space, which pushed for the use of phase space techniques for analyzing the critical properties of quantum-many body systems [73]. Recently, it has been established that the Wigner function is a bonafide measure of first-, second-, and infinite-order quantum phase transitions in the Ising and Heisenberg chains [74, 75]. In this paper, we build on these studies and we focus on the role of the Wigner function and its negative part in delimiting the phase diagram of an exotic quantum spin chain, i.e. the homogeneous spin$-(\frac{1}{2\!}-\!\frac{1}{2})$ and inhomogeneous spin$-(\frac{1}{2}\!-\!1)$ Ising-Heisenberg chains. The quasi-probability nature of the Wigner function in phase space, allows us to identify classical and quantum regimes, offering a unique perspective on quantum states and their critical properties, which is pivotal for the development of quantum technologies [76].
The relevance of our study lies in its comparison with other quantum information tools, notably the concurrence which is a measure of entanglement and has been instrumental in detecting phase boundaries in quantum systems [77, 78]. Our results highlight the advantages and limitations of using phase space methods in contrast to the entanglement concurrence. This comparison is vital in understanding the most effective approaches for studying quantum many-body systems, as different methods can yield varying insights into the critical properties of these systems. Furthermore, we shed light on the concept of the equal angle slice approximation in phase space. We show its effectiveness in capturing the essential features of the phase diagram, particularly for the homogeneous spin$-(\frac{1}{2\!}-\!\frac{1}{2})$ Ising-Heisenberg chain. Our investigation into the Wigner function’s negativity provides a profound understanding of the quantum states in these chains. In contrast, for the inhomogeneous spin$-(\frac{1}{2\!}-\!1)$ chain, we find that a more comprehensive approach, involving an integral over the entire phase space, is necessary. This distinction underscores the sensitivity of phase space methods to the homogeneity of the quantum system under consideration, highlighting the need for tailored approaches in studying different quantum systems.
@@ II The asymmetric tetrahedron Ising-Heisenberg chain
(a) Spin$-(\frac{1}{2}\!-\!\frac{1}{2})$
(b) Spin$-(\frac{1}{2}\!-\!1)$
Figure 2: Phase diagram for the asymmetric tetrahedral Ising-Heisenberg chain, Eq. (II), for two different spin configurations. Panel (a) shows the phase diagram for the spin$-(\frac{1}{2}\!-\!\frac{1}{2})$ case, utilizing parameters from Eq. (2). The spin alignments are represented by arrows, with up or down arrows indicating the two possible spin states at the Ising and Heisenberg nodes. Panel (b) illustrates the phase diagram for the spin$-(\frac{1}{2}\!-\!1)$ case, using parameters from Eq. (3). Here, up and down arrows denote the Ising nodes, while a combination of up, dot, and down illustrates the spin-1 configurations on the Heisenberg edges. Magnetic phases are labeled as FM (ferromagnetic), QFO (quantum ferromagnetic), FRI (ferrimagnetic), and FRU (frustrated ferromagnetism). The FM state is a product state, while the QFO states are entangled and differ due to the distinct orientation probabilities of spins in the Heisenberg edges. The FRI state emerge from the superposition of ferromagnetic and antiferromagnetic state, causing a competition in the spin alignment. The FRU states exhibit non-classical behaviour caused by competing interactions, leading to frustration where classical ordering is absent, suggesting a higher degree of quantum correlation. Appendix (A) elaborates on the model’s spectrum and the specific states of each phase for both scenarios.
@@
to draw the phase diagram represented in Fig. (2b). This re-parametrization essentially maps the original interaction parameters $J$, $J_{z}$, $J_{+}$, and $J_{-}$ into a new coordinate system defined by $x$ and $y$, facilitating a more tractable exploration of the model’s behavior across different parameter regimes. The mapping reveals critical points where multiple phases converge and provides a clearer understanding of the phase transitions within the ATIH chain [30].
@@ III Figures of Merit
with $d$ being the dimension of the single cell’s Hilbert space, $\sigma_{i}^{0}$ denoting the identity matrix and $\sigma_{i}^{1}=\sigma^{z}$ at site $i$ (and site $i+1$). In the same fashion, $S$ represent the spin operator with $S^{0}=\mathbb{1},S^{1}=S^{x},S^{2}=S^{y},$ and $S^{3}=S^{z}$. For the homogeneous case, i.e. spin(12,$-(\frac{1}{2},\frac{1}{2})$, $d=16$ and for the spin$-(\frac{1}{2},1)$ inhomogeneous chain $d=36$. The averages in the spin-spin correlation functions is taken over the ground state of the ATIH chain, Eq. (II), and they can be calculated using the transfer matrix approach outlined in Appendix (B). The detailed expression of the density matrix, Eq. (4), is provided in Appendix (C), from which we can compute various information-theoretic quantities, such as the entanglement measures and the Wigner function.
@@
For an $N$-partite quantum state $\rho$, residing in a composite Hilbert space represented by multiple tensor products of $\mathcal{H}$, i.e. $\rho\in\mathcal{H}\otimes\mathcal{H}\otimes...\otimes\mathcal{H}$, a foundational lower bound for the concurrence has been established [80, 81]. It is expressed as
@@
where $\hat{D}(\Omega)\hat{\Pi}\hat{D}^{\dagger}(\Omega)=\hat{\Delta}(\Omega)$ forms the kernel of this function, $\hat{\rho}$ is the density matrix describing the system and $\Omega$ represents a complete parametrization of the phase space such that $\hat{D}$ and $\hat{\Pi}$ are defined in terms of coherent states $\hat{D}(\Omega)\ket{0}=\ket{\Omega}$ and $\hat{\Pi}\ket{\Omega}=-\ket{\Omega}$. A distribution $W_{\hat{\rho}}(\Omega)$ can describe a Wigner function over a phase space parametrized by a set of $\Omega$’s, if there exists a kernel $\hat{\Delta}(\Omega)$ that generates $W_{\hat{\rho}}(\Omega)$ according to the Weyl rule
@@
[…]
@@
where $d\!=\!2j+1$ is the dimension of the Hilbert space with $j$ being the spin number, $\mathcal{N}(d)\!=\!\sqrt{d(d+1)(d-1)/2}$, and $\hat{\zeta}$ is a diagonal matrix with entries $\sqrt{\frac{2}{d(d-1)}}$ except the last element $\hat{\zeta}_{d,d}=-\sqrt{\frac{2(d-1)}{d}}$. In this case, the operator $\hat{U}(\Omega)$ represents the SU(2) rotations under three angles $\theta,\varphi,$ and $\phi$:
@@
[…]
where $d\Omega=\frac{1}{\pi}\sin(2\theta)d\theta d\phi$. This negativity has a profound physical interpretation as it is interpreted as an indicator of non-classical behavior, distinguishing quantum phenomena from their classical counterparts. In the context of quantum entanglement and quantum correlations, the negativity of the Wigner function is a valuable tool for detecting and quantifying some entangled states [85, 86]. In what follows, Equations 8, 5 and 15 will be our main figures of merit.
@@ IV Results and discussions
Figure 3: Figures of merit in the single cell spin$-(\frac{1}{2}-\frac{1}{2})$ ATIH model, Eq. (II), under the parameters defined in Eq.(2). (a) The average Wigner function values, Eq. (8), across the phase space mapping out the complete phase diagram and critical phase boundaries under the equal angle slice approximation. Distinct Wigner function values characterize different phases, varying from positive to negative as influenced by the phase space parameters $(\theta,\varphi)$, demonstrating the utility of the Wigner function in identifying phase transitions within the model. (b) The negativity of the Wigner function, Eq. (15), under the equal angle slice approximation showing its ineffectiveness in this limit in capturing properly the phase diagram (2a). We assess this properly in (c) by adopting a comprehensive mapping in the entire phase space using Monte Carlo (MC) integration, which confirms the predominance of the negative values of the Wigner function across the phase space, especially significant around the phase boundaries, except in the coherence-free FM and FRI phases. (d) Lower bound concurrence, Eq. (5), splitting the phase diagram in three parts: unentangled region which describes the FM and FRI phases. Maximally entangled region describing the QFO III and IV phases, and an intermediate region that identify the FRU III and FRU IV phases.
We begin by examining the phase diagram for the spin$-(\frac{1}{2}-\frac{1}{2})$ ATIH chain, Eq. (II). In Fig. (3a), we present the average value of Wigner function, Eq. (13), across the phase space spanned by $(\theta,\varphi)$ under the equal angle slice approximation. The Wigner function can effectively capture the entire phase diagram and all the phase boundaries (c.f Fig. (2a)). Furthermore, each phase within the diagram is characterized by distinct values of the Wigner function, which vary from positive to negative based on the parameters $(x,y)$. Here, the equal angle slice approximation is a good simplification for capturing the salient critical features of the spin$-(\frac{1}{2}-\frac{1}{2})$ ATIH model, Eq. (II), without the need to explore all the phase space which is computationally exhaustive.
@@
The negativity of the Wigner function, Eq. (15) is depicted in Fig. (3b) under the equal angle slice approximation. The negativity recognize most of the phases with distinctive values, while missing three phase boundaries, i.e. the FRI-$\text{FRU}_{III}$, the $\text{FRU}_{III}-\text{QFO}_{III}$, and the FM-QFO III phase boundaries, which are captured by the positive part of the Wigner function as shown in panel (3a). Here, the equal angle slice approximation is not adequate to properly quantify the negativity of the Wigner function in the system as we would expect all the phases of matter, except the FM and FRI phases, to have negative Wigner functions. Our expectation comes from the fact that in these phases the Heisenberg nodes responsible for the “quantum” effects in the model are described by entangled states (c.f. Equations 21, 22, 23 and 24), which are known to have a negative Wigner function [87]. We confirm this in panel (3c) where we drop the equal angle slice and compute the negativity of the Wigner function, Eq. (15), over the entire phase space spanned by the angles $\{\theta_{1},\varphi_{1};\theta_{2},\varphi_{2};\theta_{3},\varphi_{3};\theta_{4},\varphi_{4}\}$. We see that the Wigner function is negative over the entire phase diagram except in the coherence-free phases, i.e. the FM and FRI phases. Furthermore, the phase boundaries are more pronounced, including the $x\!=\!\pi$ continuous phase transition line, by a maximum amount of negativity.
@@
Figure 4: Figures of merit in the single cell spin$-(\frac{1}{2}-1)$ ATIH model, Eq. (II), under the parameters defined in Eq.(3). (a) Average of the Wigner function, Eq. (8), under the equal slice approximation showing its limitation in revealing the phase diagram (2b) for the system, which is due to the bias introduced by the approximation and the information loss in non-uniform regions of phase space. (b) The average Wigner function, Eq. (8), across the entire phase space, providing a clearer description of phase boundaries, particularly between regions of positive and negative Wigner function values. (c+d) Negativity of the Wigner function, Eq. (15), calculated via Monte Carlo integration and the lower bound concurrence, Eq. (5), respectively. Negative values cluster in phases where Heisenberg nodes align with highly entangled Bell states, such as the QFO I and QFO II states. In contrast, the FRU and QFO III phases, associated with generalized-Bell states, show no such negativity despite also being entangled. This suggests that the absence of negativity does not imply the lack of entanglement.
@@
We shift our focus to the inhomogeneous spin$-(\frac{1}{2}\!-\!1)$ chain. Similarly with the homogeneous case, we employ the equal angle slice approximation to explore the phase diagram of the system (c.f. Fig (2b)). Interestingly, this approximation is limited in this case as shown in Fig (4a), where the Wigner function fails to accurately describe the critical properties of the system. The equal angle slice method, like any approximation, sacrifices precision for computational ease. For inhomogeneous quantum systems, this can lead to issues including loss of information in non-uniform areas of the phase space, rendering the results dependent on the chosen slices. Consequently, the approximation does not capture the complete features and detailed structure of the Wigner function.
@@
Fig. (4c) and (4d) highlights the negative regions of the Wigner function across the entire phase space using the Monte Carlo approximation for computing integrals and the lower bound concurrence, respectively, for the spin$-(\frac{1}{2}-1)$ ATIH model. Negative values are localized in phases of matter where the Heisenberg nodes correspond to Bell states, specifically the QFO I and QFO II states (see Equations 53 and 54), which exhibit high entanglement as shown in panel (4d). However, the negativity of the Wigner function fail to distinguish between the FRU and QFO III phases, which involve generalized-Bell states (see Equations 58 and 55) and are also entangled according to panel (4d). This suggests that the lack of negative regions in the Wigner function does not necessarily indicate an absence of entanglement in those areas of phase space. These results call for a profound analysis of Wigner negativity and entanglement in high-dimensional states. The role of Wigner negativity in quantum systems remains a vibrant open question, particularly in exploring its potential to measure entanglement and other quantum informational resources. The measurement and practical utilization of Wigner negativity can have implications for quantum computing and quantum communication protocols.

### v2

@@ Abstract
We demonstrate the unique capabilities of the Wigner function, particularly in its positive and negative parts, for exploring the phase diagram of the spin(1212)-($\frac$$\frac{1}{2}$) and spin(121)-($\frac${1}{2}\!-\!1) Ising-Heisenberg chains. We highlight the advantages and limitations of the phase space approach in comparison with the entanglement concurrence in detecting phase boundaries. We establish that the equal angle slice approximation in the phase space is an effective method for capturing the essential features of the phase diagram, but falls short in accurately assessing the negativity of the Wigner function for the homogeneous spin(1212)-($\frac$$\frac{1}{2}$) Ising-Heisenberg chain. In contrast, we find for the inhomogeneous spin(121)-($\frac${1}{2}\!-\!1) chain that an integral over the entire phase space is necessary to accurately capture the phase diagram of the system. This distinction underscores the sensitivity of phase space methods to the homogeneity of the quantum system under consideration.
@@ I Introduction
The rapid development of quantum technologies in the 21st century allowed for the possibility of experimental measurement of the phase space, which pushed for the use of phase space techniques for analyzing the critical properties of quantum-many body systems [73]. Recently, it has been established that the Wigner function is a bonafide measure of first-, second-, and infinite-order quantum phase transitions in the Ising and Heisenberg chains [74, 75]. In this paper, we build on these studies and we focus on the role of the Wigner function and its negative part in delimiting the phase diagram of an exotic quantum spin chain, i.e. the homogeneous spin(1212)-($\frac$$\frac{1}{2}$) and inhomogeneous spin(121)-($\frac${1}{2}\!-\!1) Ising-Heisenberg chains. The quasi-probability nature of the Wigner function in phase space, allows us to identify classical and quantum regimes, offering a unique perspective on quantum states and their critical properties, which is pivotal for the development of quantum technologies [76].
The relevance of our study lies in its comparison with other quantum information tools, notably the concurrence which is a measure of entanglement and has been instrumental in detecting phase boundaries in quantum systems [77, 78]. Our results highlight the advantages and limitations of using phase space methods in contrast to the entanglement concurrence. This comparison is vital in understanding the most effective approaches for studying quantum many-body systems, as different methods can yield varying insights into the critical properties of these systems. Furthermore, we shed light on the concept of the equal angle slice approximation in phase space. We show its effectiveness in capturing the essential features of the phase diagram, particularly for the homogeneous spin(1212)-($\frac$$\frac{1}{2}$) Ising-Heisenberg chain. Our investigation into the Wigner function’s negativity provides a profound understanding of the quantum states in these chains. In contrast, for the inhomogeneous spin(121)-($\frac${1}{2\!}-\!1) chain, we find that a more comprehensive approach, involving an integral over the entire phase space, is necessary. This distinction underscores the sensitivity of phase space methods to the homogeneity of the quantum system under consideration, highlighting the need for tailored approaches in studying different quantum systems.
@@ II The asymmetric tetrahedron Ising-Heisenberg chain
(a) Spin(1212)-($\frac$$\frac{1}{2}$)
(b) Spin(121)-($\frac${1}{2}\!-\!1)
Figure 2: Phase diagram for the asymmetric tetrahedral Ising-Heisenberg chain, Eq. (II), for two different spin configurations. Panel (a) shows the phase diagram for the spin(1212)-($\frac$$\frac{1}{2}$) case, utilizing parameters from Eq. (2). The spin alignments are represented by arrows, with up or down arrows indicating the two possible spin states at the Ising and Heisenberg nodes. Panel (b) illustrates the phase diagram for the spin(121)-($\frac${1}{2}\!-\!1) case, using parameters from Eq. (3). Here, up and down arrows denote the Ising nodes, while a combination of up, dot, and down illustrates the spin-1 configurations on the Heisenberg edges. Magnetic phases are labeled as FM (ferromagnetic), QFO (quantum ferromagnetic), FRI (ferrimagnetic), and FRU (frustrated ferromagnetism). The FM state is a product state, while the QFO states are entangled and differ due to the distinct orientation probabilities of spins in the Heisenberg edges. The FRI state emerge from the superposition of ferromagnetic and antiferromagnetic state, causing a competition in the spin alignment. The FRU states exhibit non-classical behaviour caused by competing interactions, leading to frustration where classical ordering is absent, suggesting a higher degree of quantum correlation. Appendix (A) elaborates on the model’s spectrum and the specific states of each phase for both scenarios.
@@
to draw the phase diagram represented in Fig. (2b). This re-parametrization essentially maps the original interaction parameters $J$, $_{z}, J+J_{+}$, and $J_{-}$ into a new coordinate system defined by $x$ and $y$, facilitating a more tractable exploration of the model’s behavior across different parameter regimes. The mapping reveals critical points where multiple phases converge and provides a clearer understanding of the phase transitions within the ATIH chain [30].
@@ III Figures of Merit
with $d$ being the dimension of the single cell’s Hilbert space, $\sigma_{i}^{0}$ denoting the identity matrix and $\sigma_{i}^{1}=\sigma^{z}$ at site $i$ (and site $i+1$). In the same fashion, $S$ represent the spin operator with $S^{0}=\mathbb{1},S^{1}=S^{x},S^{2}=S^{y},$ and $S^{3}=S^{z}$. For the homogeneous case, i.e. spin(12,12)-($\frac$$\frac{1}{2}$), $d=16$ and for the spin(12,1)-($\frac${1}{2},1) inhomogeneous chain $d=36$. The averages in the spin-spin correlation functions is taken over the ground state of the ATIH chain, Eq. (II), and they can be calculated using the transfer matrix approach outlined in Appendix (B). The detailed expression of the density matrix, Eq. (4), is provided in Appendix (C), from which we can compute various information-theoretic quantities, such as the entanglement measures and the Wigner function.
@@
For an $N$-partite quantum state $\rho$, residing in a composite Hilbert space represented by multiple tensor products of $\mathcal{H}$, i.e. $\rho\in\mathcal{H}\otimes\mathcal{H}\otimes$$\otimes\mathcal{H}$, a foundational lower bound for the concurrence has been established [80, 81]. It is expressed as
@@
[…]
@@
and, as stated in [84], also satisfy the Stratonovich-Weyl correspondences, which articulate several foundational properties of the Wigner function. Primarily, they allow for a bi-directional reconstruction between the density matrix $\hat{\rho}$ and its representation $W_{\hat{\rho}}(\Omega)$ through a precise mathematical mapping. This mapping not only facilitates the transition from $\hat{\rho}$ to $W_{\hat{\rho}}(\Omega)$ via the trace formula $W_{\hat{\rho}}(\Omega)=\text{Tr}\left(\hat{\rho}\hat{\Delta}(\Omega)\right)$ but also enables the reconstruction of $\hat{\rho}$ from $W_{\hat{\rho}}(\Omega)$ by integrating over $\Omega$ with the kernel $\hat{\Delta}(\Omega)$. Another critical aspect of these correspondences is the reality and normalization to unity of $W_{\hat{\rho}}$, ensuring that it always represents a valid quantum state. Additionally, the invariance of $\hat{\rho}$ under global unitary operations implies a corresponding invariance in $W_{\hat{\rho}}$, preserving the physical characteristics of the quantum state within the phase space. Lastly, a distinctive feature of the Wigner function is its ability to quantify the overlap between two states $\hat{\rho}^{\prime}$ and $\hat{\rho}^{\prime\prime}$. This is executed through a definite integral over $\Omega$ via $\int_{\Omega}W_{\hat{\rho}^{\prime}}W_{\hat{\rho}^{\prime\prime}}d\Omega=\text{Tr}\left(\hat{\rho}^{\prime}\hat{\rho}^{\prime\prime}\right)$ which highlights a unique property that sets the Wigner function apart in the analysis of quantum states.
An extension of Eq. (8) to finite-dimensional systems requires the construction of a kernel $\hat{\Delta}(\Omega)$ that reflects the symmetries of the system at hand. For spin systems, Tilma et al [84] argued that their Wigner functions can be generated via a spinj-j representation of SU(2) under the following kernels:
@@
where d=2j+1d\!=\!2j+1 is the dimension of the Hilbert space with $j$ being the spin number, $\mathcal{N}(d)\!=\!\sqrt{d(d+1)(d-1)/2}$, and $\hat{\zeta}$ is a diagonal matrix with entries $\sqrt{\frac{2}{d(d-1)}}$ except the last element $\hat{\zeta}_{d,d}=-\sqrt{\frac{2(d-1)}{d}}$. In this case, the operator $\hat{U}(\Omega)$ represents the SU(2) rotations under three angles $\theta,\varphi,$ and $\phi$:
@@
$\mathcal{N}_{W}=\frac{1}{2}\int_{\Omega}\Big(|W_{\rho}(\Omega)|-W_{\rho}(\Omega)\Big)d\Omega.$ (15)
where $d\Omega=\frac{1}{\pi}\sin(2\theta)d\theta d\phi$. This negativity has a profound physical interpretation as it is interpreted as an indicator of non-classical behavior, distinguishing quantum phenomena from their classical counterparts. In the context of quantum entanglement and quantum correlations, the negativity of the Wigner function is a valuable tool for detecting and quantifying some entangled states [85, 86]. In what follows, Equations˜8, 5 and 15 will be our main figures of merit.
@@ IV Results and discussions
Figure 3: Figures of merit in the single cell spin(1212)-($\frac$$\frac{1}{2}$) ATIH model, Eq. (II), under the parameters defined in Eq.(2). (a) The average Wigner function values, Eq. (8), across the phase space mapping out the complete phase diagram and critical phase boundaries under the equal angle slice approximation. Distinct Wigner function values characterize different phases, varying from positive to negative as influenced by the phase space parameters $(\theta,\varphi)$, demonstrating the utility of the Wigner function in identifying phase transitions within the model. (b) The negativity of the Wigner function, Eq. (15), under the equal angle slice approximation showing its ineffectiveness in this limit in capturing properly the phase diagram (2a). We assess this properly in (c) by adopting a comprehensive mapping in the entire phase space using Monte Carlo (MC) integration, which confirms the predominance of the negative values of the Wigner function across the phase space, especially significant around the phase boundaries, except in the coherence-free FM and FRI phases. (d) Lower bound concurrence, Eq. (5), splitting the phase diagram in three parts: unentangled region which describes the FM and FRI phases. Maximally entangled region describing the QFO III and IV phases, and an intermediate region that identify the FRU III and FRU IV phases.
We begin by examining the phase diagram for the spin(1212)-($\frac$$\frac{1}{2}$) ATIH chain, Eq. (II). In Fig. (3a), we present the average value of Wigner function, Eq. (13), across the phase space spanned by $(\theta,\varphi)$ under the equal angle slice approximation. The Wigner function can effectively capture the entire phase diagram and all the phase boundaries (c.f Fig. (2a)). Furthermore, each phase within the diagram is characterized by distinct values of the Wigner function, which vary from positive to negative based on the parameters $(x,y)$. Here, the equal angle slice approximation is a good simplification for capturing the salient critical features of the spin(1212)-($\frac$$\frac{1}{2}$) ATIH model, Eq. (II), without the need to explore all the phase space which is computationally exhaustive.
@@
The negativity of the Wigner function, Eq. (15) is depicted in Fig. (3b) under the equal angle slice approximation. The negativity recognize most of the phases with distinctive values, while missing three phase boundaries, i.e. the FRI-$\text{FRU}_{III}$, the $\text{FRU}_{III}-\text{QFO}_{III}$, and the FM-QFO III phase boundaries, which are captured by the positive part of the Wigner function as shown in panel (3a). Here, the equal angle slice approximation is not adequate to properly quantify the negativity of the Wigner function in the system as we would expect all the phases of matter, except the FM and FRI phases, to have negative Wigner functions. Our expectation comes from the fact that in these phases the Heisenberg nodes responsible for the “quantum” effects in the model are described by entangled states (c.f. Equations˜21, 22, 23 and 24), which are known to have a negative Wigner function [87]. We confirm this in panel (3c) where we drop the equal angle slice and compute the negativity of the Wigner function, Eq. (15), over the entire phase space spanned by the angles $\{\theta_{1},\varphi_{1};\theta_{2},\varphi_{2};\theta_{3},\varphi_{3};\theta_{4},\varphi_{4}\}$. We see that the Wigner function is negative over the entire phase diagram except in the coherence-free phases, i.e. the FM and FRI phases. Furthermore, the phase boundaries are more pronounced, including the $x\!=\!\pi$ continuous phase transition line, by a maximum amount of negativity.
@@
Figure 4: Figures of merit in the single cell spin(121)-($\frac${1}{2}-1) ATIH model, Eq. (II), under the parameters defined in Eq.(3). (a) Average of the Wigner function, Eq. (8), under the equal slice approximation showing its limitation in revealing the phase diagram (2b) for the system, which is due to the bias introduced by the approximation and the information loss in non-uniform regions of phase space. (b) The average Wigner function, Eq. (8), across the entire phase space, providing a clearer description of phase boundaries, particularly between regions of positive and negative Wigner function values. (c+d) Negativity of the Wigner function, Eq. (15), calculated via Monte Carlo integration and the lower bound concurrence, Eq. (5), respectively. Negative values cluster in phases where Heisenberg nodes align with highly entangled Bell states, such as the QFO I and QFO II states. In contrast, the FRU and QFO III phases, associated with generalized-Bell states, show no such negativity despite also being entangled. This suggests that the absence of negativity does not imply the lack of entanglement.
@@
We shift our focus to the inhomogeneous spin(121)-($\frac${1}{2}\!-\!1) chain. Similarly with the homogeneous case, we employ the equal angle slice approximation to explore the phase diagram of the system (c.f. Fig (2b)). Interestingly, this approximation is limited in this case as shown in Fig (4a), where the Wigner function fails to accurately describe the critical properties of the system. The equal angle slice method, like any approximation, sacrifices precision for computational ease. For inhomogeneous quantum systems, this can lead to issues including loss of information in non-uniform areas of the phase space, rendering the results dependent on the chosen slices. Consequently, the approximation does not capture the complete features and detailed structure of the Wigner function.
@@
Fig. (4c) and (4d) highlights the negative regions of the Wigner function across the entire phase space using the Monte Carlo approximation for computing integrals and the lower bound concurrence, respectively, for the spin(121)-($\frac${1}{2}-1) ATIH model. Negative values are localized in phases of matter where the Heisenberg nodes correspond to Bell states, specifically the QFO I and QFO II states (see Equations˜53 and 54), which exhibit high entanglement as shown in panel (4d). However, the negativity of the Wigner function fail to distinguish between the FRU and QFO III phases, which involve generalized-Bell states (see Equations˜58 and 55) and are also entangled according to panel (4d). This suggests that the lack of negative regions in the Wigner function does not necessarily indicate an absence of entanglement in those areas of phase space. These results call for a profound analysis of Wigner negativity and entanglement in high-dimensional states. The role of Wigner negativity in quantum systems remains a vibrant open question, particularly in exploring its potential to measure entanglement and other quantum informational resources. The measurement and practical utilization of Wigner negativity can have implications for quantum computing and quantum communication protocols.